F=20N,m1=m2=3and acceleration is 0.5m/s^2.
Friction forces on the two blocks are eaqual what is magnitude

Grade:10

1 Answers

Arun
5 years ago
First consider the 2 masses as a single object. Total mass = 3+3=6 kg 
Resultant force, F= ma = 6 x 0.5 = 3N 
Since an external force of 20N is being applied, the frictional force must be 20-3=17N 
(8.5N on each mass). 
Now consider m1 by itself. The resultant force on it = ma = 3 x 0.5=1.5N 
Let T = tension in rope joining m1 and m2 
The only 2 horizontal forces on m1 are T (to right) and 8.5N (friction to the left) 
Resultant horizontal force on m1 = T1 - 8.5 
T-8.5 = 1.5 
T = 10N